a boat can travel on a river 4 miles upstream in the same time it takes to travel 7 miles downstream. find the speed of the current if the speed of the boat in still water is 11 miles per hour
Answer by stanbon(75887)

a boat can travel on a river 4 miles upstream in the same time it takes to travel 7 miles downstream. find the speed of the current if the speed of the boat in still water is 11 miles per hour
------
Upstream DATA:
distance = 4 miles ; rate = 11-c ; time = 4/(11-c)
-----------------
Downstream DATA:
distance = 7 miles ; rate = 11+c ; time = 7/(11+c)
-------
Equation:
time = time
4/(11-c) = 7/(11+c)
4(11+c) = 7(11-c)
44 + 4c = 77-7c
11c = 33
c = 3 mph (speed of the current)
=====================
Cheers,
Stan H.
-----------------